OpenStudy (anonymous):

w2+5w+4 factor the trinominal

OpenStudy (anonymous):

(w+4)(w+1)

OpenStudy (anonymous):

w = -4, w = -1

OpenStudy (anonymous):

\[w^2+5w+4\] \[w^2+4w+1w+4\] \[w(w+4)+1(w+4)\] \[(w+1)(w+4)\] either w+1=0 or w+4=0 hence w=-1 or w=-4

OpenStudy (anonymous):

got it nemokimo ? ?

OpenStudy (anonymous):

well since it is not given that \[w^2+5w+4=0\] hence we can not say that w=-1 or w=-4....